The Square Foot Grill, Inc. issued $214,000 of 10-year, 5 percent bonds on January 1, 2018, at 102. interest is payable in cash annually on December 31. The straight-line method is used for amortization.

A. Determine the carrying value (face value less discount or plus premium) of the bond liability as of December 31, 2018.
B. Determine the amount of interest expense reported on the 2018 income statement.
C. Determine the carrying value of the bond liability as of December 31, 2019.
D. Determine the amount of interest expense reported on the 2019 income statement.
Answer #1

Par value of bonds = $214,000

Issue price of bonds = 214,000 x 102%

= $218,280

Premium on bonds payable = Issue price of bonds - Par value of bonds  

= 218,280 - 214,000

= $4,280

Premium on bonds payable to be amortized annually = 4,280/10

= $428

Annual interest on bonds = 214,000 x 5%

= $10,700

Annual interest expense = Annual interest on bonds - Premium on bonds payable to be amortized annually

= 10,700 - 428

= $10,272

(A) Carrying value of bonds on Dec. 31, 2018 = Premium on bonds payable - Bonds premium amortized in 2018

= 218,280 - 428

= $217,852

(B) Interest expense reported on the 2018 income statement = $10,272

(C) carrying value of the bond liability as of December 31, 2019 = Carrying value of bonds on Dec. 31, 2018 - Bonds premium amortized in 2018

= 217,852 - 428

= $217,424

(D) Interest expense reported on the 2019 income statement = $10,272

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.